OpenStudy (anonymous):

What statement correctly describes the key features of the graph of f(x) = −3(1/3)^x + 1 − 2?

OpenStudy (anonymous):

Y-intercept of (0, −3), starts down on the left, gets closer to y = −2 on the right Y-intercept of (0, −3), starts up on the left, gets closer to y = −2 on the right Y-intercept of (0, 2), starts down on the left, gets closer to y = −1 on the right Y-intercept of (0, 2), starts up on the left, gets closer to y = −1 on the right

OpenStudy (anonymous):

@pooja195

OpenStudy (irishboy123):

and you say ... ?!?!

OpenStudy (anonymous):

A!

OpenStudy (irishboy123):

is it: \(\large f(x) = −3(\frac{1}{3})^{x + 1} − 2\)

OpenStudy (anonymous):

yes

OpenStudy (irishboy123):

\(f(x) = -3 (\frac{1}{3})^x . (\frac{1}{3}) -2 \\ = -(\frac{1}{3})^x -2\) so \(f(0) = -3\) and \(f(\infty) = -2\) i'd have to agree with you

OpenStudy (anonymous):

@IrishBoy123 Would the graph start down on the left too?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (irishboy123):

so, if you start with \(\large f(x)=−(\frac{1}{3})^x−2\), then \(\large f(-100) = −(\frac{1}{3})^{-100}−2 \\ \large = −(\frac{3}{1})^{+100}−2 = −(3)^{+100}−2\) ie a very large negative number.....

OpenStudy (irishboy123):

\(f(-1) = -(3)^1 - 2 = -5\) play with the numbers